OpenStudy (anonymous):

I need an explanation: What is the sum of a 6–term geometric series if the first term is 6 and the last term is 1,000,842?

OpenStudy (anonymous):

?

OpenStudy (mathmate):

Divide 1000842 by 6 and take the 5th root of the quotient.

OpenStudy (mathmate):

Let p = the common ratio, then first term is a0, and the 6th term is ar^5. So a0 r^5/a0=r^5. The fifth-root of the quotient is therefore the common ratio (=7)

OpenStudy (mathmate):

correction r=common ratio.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

a(sub n)=a(sub 1)r^(n-1)

OpenStudy (mathmate):

You're welcome! But see what LivyLou has to say. She has been working on your problem for quite a while.

OpenStudy (anonymous):

Oops yeah I took awhile because I was checking to see if I was giving the right formula

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

what do i put into the formula

OpenStudy (anonymous):

i know the common ratio is 7 but how do u get that

OpenStudy (anonymous):

I'm actually not sure either. I reread the question and I don't think the formula I gave helps. sorry

OpenStudy (anonymous):

its okay

OpenStudy (mathmate):

@jittcam @LivyLou The first term can be considered as a(1), and the second term as a(2), ... the nth term is a(n). We know that there is a common ratio, r, such that \[ r=a_2 / a_1 = a_3 / a_2 = .... \] So \[ a_0*r = a_1 \] \[ a_0*r^2 = a_2 \] \[ a_0*r^3 = a_3 \] ... \[ a_0*r^6 = a_6 \] Thus \[ \frac{a_6}{a_1} = \frac{a_0 r^6}{a_0 r} = r^5 \] In our case, \[ \frac{a_6 }{a_1} = \frac{1000842}{6} = r^5 \] Therefore \[ r=166807^{1/5} = 7\] Hope this helps.

OpenStudy (anonymous):

when i put 166807^{1/5} in my calculator i get 11 not 7